OpenStudy (wentz):

A right triangle has a leg b of length 7 and a hypotenuse of length 11. What is the length of the other leg, a? Round to the nearest tenth if necessary. A. 2.0 B. 8.5 C. 13.0 D. 9.6

OpenStudy (dayakar):

can u draw a rough figure

OpenStudy (igreen):



OpenStudy (igreen):

Use the pythagorean theorem. \(\sf a^2 + b^2 = c^2\) Where 'a' and 'b' are the two legs and 'c' is the hypotenuse. We are given one of the legs and the hypotenuse. \(\sf a^2 + 7^2 = 11^2\) Solve for 'a'

OpenStudy (igreen):

@wentz